# Evaluating Morphological Compositional Generalization in Large Language Models

### v2

@@ Abstract
Mete Ismayilzada1,2, Defne Circi*3, Jonne Sälevä*4, Hale Sirin7, Abdullatif Köksal5,6, Bhuwan Dhingra3, Antoine Bosselut1, Lonneke van der Plas 2,8, Duygu Ataman9 1EPFL, 2Idiap Research Institute, 3Duke University, 4Brandeis University, 5LMU Munich, 6University of Cambridge, 7Johns Hopkins University, 8Università della Svizzera Italiana, 9New York University mete.ismayilzada@idiap.ch
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable advances in the broad domain of natural language generation and understanding tasks Gemini (2024); Zhao et al. (2023); Bubeck et al. (2023); Wei et al. (2022); Brown et al. (2020). However, these models have also been shown to lack strong linguistic generalization capabilities Weissweiler et al. (2023); McCoy et al. (2023); Goldman et al. (2022); Wilson et al. (2023); Linzen (2020); Baroni (2019). This discrepancy casts doubt on whether language models learn a language the same way as humans do. When learning a language which is essentially a finite set of words and rules, humans exhibit linguistic creativity Chomsky (1965); Bergs (2019) and compositional generalization by producing and understanding novel combinations of these grammar units which are respectively known as the productivity and systematicity aspects of compositionality Fodor and Pylyshyn (1988); Chomsky (1957). While compositional generalization abilities of language models have been extensively studied Lake and Baroni (2018); Keysers et al. (2019); Kim and Linzen (2020), the extent to which language models employ this ability in morphology however remains largely under-explored. Recent works evaluating morphological generalization in language models have only focused on the productivity aspect with a limited coverage of inflectional forms Weissweiler et al. (2023); Anh et al. (2024).
@@
In summary, our contributions are as follows: 1) We design novel morphological generalization tasks that require compositional processing. 2) We prepare specific test suites in both Turkish and Finnish to measure morphological generalization and make these available for future research. 3) Using our novel tasks and test suites, we conduct a systematic analysis of morphological compositional generalization abilities of LLMs. 4) Our findings reveal a systematic gap in LLM’s ability compared to humans concerning morphological generalization that also requires compositionality.
@@ 2.1 Compositional Generalization
Compositional generalization is the capacity to understand and produce novel compositions of seen primitives and is typically characterized by systematicity and productivity Fodor and Pylyshyn (1988); Keysers et al. (2019). Systematicity refers to the ability to understand different combinations that are made up of the same known components such as John loves Mary and Mary loves John. Productivity, on the other hand, is the ability to produce potentially infinite novel combinations of a finite number of known building blocks such as using conjunctions to construct sentences Mary knows that John loves Mary and John heard that Mary knows that John loves Mary. Past work has developed several benchmarks to measure compositional generalization abilities of neural models both in fine-tuning and in-context learning settings and has shown this task to be highly challenging Yang et al. (2024); Lake and Baroni (2018); Keysers et al. (2019); Kim and Linzen (2020); An et al. (2023); Dziri et al. (2023). These benchmarks have mainly focused on synthetic sequence matching, semantic parsing, question-answering and problem-solving tasks. Our work however, investigates compositional generalization in the context of morphology.
@@ 2.2 Morphological Generalization
Morphological generalization is the ability to understand words based on their constituent parts known as morphemes and combine them to derive new words Wysocki and Jenkins (1987). Morphemes are the smallest meaningful units of language that typically correspond to word roots and affixes (i.e. prefixes, infixes and suffixes). Composing these units to construct new words can be done through inflection and derivation tasks in morphology where derivation changes the syntactic category of the words and inflection does not. These tasks have gained considerable attention as part of the SIGMORPHON’s shared tasks Cotterell et al. (2016, 2018); Vylomova et al. (2020); Kodner and Khalifa (2022); Goldman et al. (2023) and efforts to create a universal morphology McCarthy et al. (2020). While transformer-based models have been shown to achieve near-perfect accuracy on these tasks Canby et al. (2020), recent work has also found that these results are inflated due to lemma overlap pointing to a lack of generalization Goldman et al. (2022). Other works have recently investigated the morphological capabilities of LLMs using inflection tasks and reported similarly weak performance results Anh et al. (2024); Weissweiler et al. (2023). Similar to our study, both of these works use the popular Wug test Berko Gleason (1958) to evaluate the morphological generalization, however, they only focus on the productivity aspect, and their coverage of inflectional and derivational forms is limited. On the other hand, we cast the inflection and derivation tasks into the form of a compositional generalization task and evaluate models on both productivity and systematicity aspects. While focus of other works is breadth (languages from different families), we instead conduct an in-depth analysis of morphological generalization in typologically similar but unrelated languages with a large test suite covering a wide and diverse range of inflectional and derivational combinations.
@@ 3.1 Tasks
We design two novel and simple compositional tasks to test morphological abilities of models. First, a morphological productivity task which we define as a generative task where the model is given a word root, a list of affixes (not necessarily in the correct order) and is asked to derive a meaningful word by composing the root with the affixes in the correct order. Second, a morphological systematicity task which we define as a binary discriminative task where the model is again given a word root, a list of affixes and a word derived from the root using the given affixes (not necessarily a meaningful word) and is asked to determine the grammatical validity of the derived word. Figure 1 illustrates these tasks with an example in Turkish.
@@ 3.2 Data
We focus our study on two highly agglutinative languages, Turkish and Finnish, and prepare test suites  [^2] Link to the test suites will be available in camera-ready version. specific for our tasks in these languages. We particularly choose these languages because they are characterized by a large number of morphemes and hence require a high degree of compositional generalization ability.
@@
We first collect a $\sim$1,000,000 sentence subsample of the Finnish mC4 corpus (Xue et al., 2021). We then extract unique words from the text and morphologically segment them using omorfi (Pirinen, 2015) and UralicNLP (Hämäläinen, 2019). After excluding words that analyzers did not cover, we manually annotate the segmentations to identify prefixes, lemmas, and affixes among the segments. We then perform stratified sampling based on the number of affixes to ensure an even range of morphological complexity in our data set. Finally, we extract sentences corresponding to each analyzed word from mC4 and validate whether they make sense. In a significant portion of cases, we notice that the raw sentences are noisy; in these cases, we opt to generate synthetic sentences using ChatGPT, which we then validate to be grammatical. Final data statistics and examples can be found in Appendix Tables 6 and 7.
@@ 4 Experiments
We treat the productivity task as an open-ended task in which the model is asked to derive a word from the given root and affixes and the systematicity task as a binary classification task in which the model is asked to determine whether the given derivation is grammatically correct. For the systematicity task, we generate negative examples by producing all the combinations of morphemes attached to the same root and choosing the top four compositions (two for morpheme lengths of 1 and 2) [^4] For 1-morpheme words, we manually annotate a negative morpheme to generate one negative option. that are closest to the original valid combination measured by the Levenshtein distance. We do this to ensure our incorrect combinations are challenging enough for the model as they will be deceptively close to a plausible derivation. We also experiment with other negative example selection strategies such as random selection and a heuristic selection based on the linguistic characteristics of the given language. We describe these settings in more detail and compare the results in Section 5.5. Finally, we manually verify all the generated negative examples and fix the label of false negatives.
@@
We evaluate several state-of-the-art multilingual instruction-finetuned LLMs, namely, two open-weights models, Aya-23 Aryabumi et al. (2024) and Qwen-2.5 Team (2024), and two closed-source models, Gemini-1.5 Gemini (2024) and GPT-4 OpenAI (2024). We evaluate all models on all languages except for Aya-23 which officially supports Turkish, but not Finnish.  [^5] We also experimented with recent LLMs that are instruction-finetuned specifically on Finnish such as Poro-34B Luukkonen et al. (2024) and Ahma model series that are Llama Touvron et al. (2023) models fine-tuned on Finnish, however, we omitted them from our analysis as they failed to follow our task prompts in both English and Finnish templates.. We also report the performance of a random baseline that generates a derivation with a random combination of given morphemes (productivity task) and randomly decides whether the derivation is grammatically correct, and a majority baseline, which selects the most frequent label (in our case "No") for the systematicity task (not applicable for the productivity task). All models are evaluated using few-shot (1, 3, and 5) in-context learning and greedy decoding since our tasks are deterministic by nature. Unless otherwise specified, prompt instructions are in English, and number of shots is set to $5$ for reported results. Further details on model evaluation can be found in Appendix C.
@@
For the productivity task, we use Exact Match accuracy against the correct derivations. For the systematicity task, we report an average of Macro-F1 scores for each sample and a Coherence score that measures whether the model correctly and consistently identifies the validity (or invalidity) of all derivations for a given set of morphemes. We employ this stringent metric to test the robustness of model performance similar to Storks and Chai (2021).
@@ 5.1 Effect of Morphological Complexity
Recent works have shown that morphological complexity plays a crucial role in the morphological generalization abilities of LLMs Anh et al. (2024). Morphological complexity is typically categorized into integrative (I-complexity) which refers to the predictability of inflected form and enumerative (E-complexity) complexity which refers to the number of cases and inflectional paradigms in language grammar Ackerman and Malouf (2013). While both languages we study are morphologically complex, our test suites include inflectional and derivational forms of varying length in the number of morphemes (1-7 in Turkish and 1-6 in Finnish). This allows us to study the effect of within-language E-complexity on the performance of our models. Figure 4 summarizes the GPT-4 performance for both tasks stratified by the number of bound morphemes on the Turkish data. On the productivity task, we observe a sharp downward trend (plummeting to nearly zero) in performance as the number of morphemes increases for both ID and OOD test suites with a relatively constant gap between ID and OOD performance while humans exhibit no such dependence on complexity (Appendix Tables 16, 25). This shows that humans learn their native language robustly and can easily produce and identify long novel words while models are quite sensitive to the morphological (E-) complexity.
@@ 5.3 Effect of Tokenization
Past work has shown that suboptimal tokenizers, especially byte-pair encoding Sennrich et al. (2016) used in GPT-4 have generally a negative effect on the morphological abilities of language models Meyer and Buys (2023); Bostrom and Durrett (2020); Hofmann et al. (2021). Whether the low performance of the model on the productivity task can be attributed to the suboptimal nature of the tokenization is of interest in particular because our tasks rely on the morphologically segmented morphemes while the model utilizes byte-level tokens that are mostly English. To measure the effect of the tokenization, we ran a version of the productivity task where the morphemes provided to the model are obtained by segmenting the final derivation based on the model’s own tokenizer instead of the morphologically-aligned units. Figure 6 compares the performance of the tokenizer-aligned morphemes with the morphologically-aligned morphemes on the ID test set. [^7] Since we use the word root as a definition for the nonce root and the tokenizer tends to break the words into meaningless chunks, we skip this experiment on the OOD test set. We see that the performance in both cases is very similar to each other which points to a possibility that tokenization may not be the underlying issue behind the low performance. This finding is also consistent with some past work on exploring morphological capabilities of ChatGPT Weissweiler et al. (2023).
@@ 6 Conclusion
In this paper, we proposed a novel experimental paradigm to test morphological generalization abilities of large language models through compositionality. Our tasks target measuring morphological productivity and systematicity in a given language. We applied these tasks on the morphologically complex languages of Turkish and Finnish and evaluated morphological compositional generalization abilities of several state-of-the-art large language models. Our experimental results and analysis reveal a significant gap in the performance of LLMs compared to humans with respect to generalization in morphology.
@@
While our novel tasks are language, dataset, and model-independent, our study only focused on two agglutinative languages and a few large language models. Therefore, the applicability of our findings in other languages and models should be further studied. We also mainly focused on the grammatical validity of the words, whereas it would be equally interesting to study the capacity of LLMs to produce and understand novel semantically and pragmatically valid derivations. While we have also optimized our prompts to be as simple and maximally instructive and tested in multiple languages and in chain-of-thought setting, whether a different set of prompts would produce the same results is not clear. Finally, we only evaluate models using greedy decoding due to the deterministic nature of our tasks, however, the effect of different decoding strategies need to be explored.

### v4

@@ Abstract
Mete Ismayilzada1,2, Defne Circi*3, Jonne Sälevä*4, Hale Sirin7, Abdullatif Köksal5,6, Bhuwan Dhingra3, Antoine Bosselut1, Duygu Ataman†9, Lonneke van der Plas †2,8 1EPFL, 2Idiap Research Institute, 3Duke University, 4Brandeis University, 5LMU Munich, 6University of Cambridge, 7Johns Hopkins University, 8Università della Svizzera Italiana, 9New York University mahammad.ismayilzada@epfl.ch
@@ 1 Introduction
Large language models (LLMs) have recently achieved remarkable advances in the broad domain of natural language generation and understanding tasks Gemini (2024); Zhao et al. (2023); Bubeck et al. (2023); Wei et al. (2022); Brown et al. (2020). However, these models have also been shown to lack strong linguistic generalization capabilities Weissweiler et al. (2023); McCoy et al. (2023); Goldman et al. (2022); Wilson et al. (2023); Linzen (2020); Baroni (2019). This discrepancy casts doubt on whether language models learn a language the same way as humans do. When learning a language which is essentially a finite set of words and rules, humans exhibit linguistic creativity Chomsky (1965); Bergs (2019) and compositional generalization through productivity and systematicity Fodor and Pylyshyn (1988a); Chomsky (1957). These abilities allow humans respectively to produce and understand novel combinations of familiar grammar units. While compositional generalization abilities of language models have been extensively studied Lake and Baroni (2018); Keysers et al. (2019); Kim and Linzen (2020), the extent to which language models employ this ability in morphology however remains largely under-explored. Recent works evaluating morphological generalization in language models have only focused on the productivity aspect with a limited coverage of inflectional forms Weissweiler et al. (2023); Anh et al. (2024).
@@
In summary, our contributions are as follows: 1) We design novel morphological generalization tasks that require compositional processing. 2) We prepare specific test suites in both Turkish and Finnish to measure morphological generalization and make these available for future research [^2] https://github.com/mismayil/morph-gen. 3) Using our novel tasks and test suites, we conduct a systematic analysis of morphological compositional generalization abilities of LLMs. 4) Our findings reveal a systematic gap in LLM’s ability compared to humans concerning morphological generalization in agglutinative languages that also requires compositionality.
@@ 2.1 Compositional Generalization
Compositional generalization is the capacity to understand and produce novel compositions of seen primitives and is typically characterized by systematicity and productivity Fodor and Pylyshyn (1988a); Keysers et al. (2019). Systematicity refers to the ability to understand different combinations that are made up of the same known components such as John loves Mary and Mary loves John. Productivity, on the other hand, is the ability to produce potentially infinite novel combinations of a finite number of known building blocks such as using conjunctions to construct sentences Mary knows that John loves Mary and John heard that Mary knows that John loves Mary. Past work has developed several benchmarks to measure compositional generalization abilities of neural models both in fine-tuning and in-context learning settings and has shown this task to be highly challenging Yang et al. (2024); Lake and Baroni (2018); Keysers et al. (2019); Kim and Linzen (2020); An et al. (2023); Dziri et al. (2023). These benchmarks have mainly focused on synthetic sequence matching, semantic parsing, question-answering and problem-solving tasks. Our work however, investigates compositional generalization in the context of morphology.
@@ 2.2 Morphological Generalization
Morphological generalization is the ability to understand words based on their constituent parts known as morphemes and combine them to derive new words Wysocki and Jenkins (1987). Morphemes are the smallest meaningful units of language that typically correspond to word roots and affixes (i.e. prefixes, infixes and suffixes). Composing these units to construct new words can be done through inflection and derivation tasks in morphology where derivation often changes the syntactic category of the words and inflection does not. These tasks have gained considerable attention as part of the SIGMORPHON’s shared tasks Cotterell et al. (2016, 2018a); Vylomova et al. (2020); Kodner and Khalifa (2022); Goldman et al. (2023) and efforts to create a universal morphology McCarthy et al. (2020). While transformer-based models have been shown to achieve near-perfect accuracy on these tasks Canby et al. (2020), recent work has also found that these results are inflated due to lemma overlap pointing to a lack of generalization Goldman et al. (2022). Other works have recently investigated the morphological capabilities of LLMs using inflection tasks and reported similarly weak performance results Anh et al. (2024); Weissweiler et al. (2023). Similar to our study, both of these works use the popular Wug test Berko Gleason (1958) to evaluate the morphological generalization, however, they only focus on the productivity aspect, and their coverage of inflectional and derivational forms is limited. For example, Weissweiler et al. (2023) considers only a handful of specific inflectional forms (e.g. first person singular agreement and past tense, second person plural agreement etc.) for each language and Anh et al. (2024) translates the original Wug test suite which is very small in size (23 samples) into different languages. On the other hand, we cast the inflection and derivation tasks into the form of a compositional generalization task and evaluate models on both productivity and systematicity aspects. While focus of other works is breadth (languages from different families), we instead conduct an in-depth analysis of morphological generalization in typologically similar but unrelated languages with a large test suite covering a wide and diverse range of inflectional and derivational combinations.
@@ 3 Methodology
3.1 Background
The important role of compositional processing in language understanding and generation has been extensively studied Carnap (1947); Chomsky (1965); Fodor and Pylyshyn (1988b); Zadrozny (1994); Bauer (2001); Aronoff and Lindsay (2014). Past works have shown that new word formation is often a multi-level process that requires identifying the correct order of primary and secondary morphemes Kiparsky (1982a, b); Hockett (1954), and while humans might memorize some frequent words and phrases as a whole, most of the expressive language generation relies on productive rules of grammar O’Donnell (2015). However, not all languages are equally productive, and more productive languages (e.g. agglutinative) tend to have complex inflectional morphology Cotterell et al. (2019); Ackerman and Malouf (2013). Moreover, these languages have been shown to be harder to model for n-gram and recurrent language models Cotterell et al. (2018b); Czarnowska et al. (2019). Inspired by these works, we focus our study on two highly agglutinative languages and compositional tasks which we describe in detail below.
@@ 3.2 Tasks
Similar to works studying compositional abilities of neural networks Goodwin et al. (2020); Lake and Baroni (2018); Keysers et al. (2019), we design two novel and simple compositional probing tasks to test morphological abilities of models. First, a morphological productivity task which we define as a generative task where the model is given a word root, a list of affixes (not necessarily in the correct order) and is asked to derive a meaningful word by composing the root with the affixes in the correct order. Second, a morphological systematicity task which we define as a binary discriminative task where the model is again given a word root, a list of affixes and a word derived from the root using the given affixes (not necessarily a meaningful word) and is asked to determine the grammatical validity of the derived word. Figure 1 illustrates these tasks with an example in Turkish.
@@ 3.3 Data
We focus our study on two highly agglutinative languages, Turkish and Finnish, and prepare test suites specific for our tasks in these languages. We particularly choose these languages because they are characterized by a large number of morphemes and hence require a high degree of compositional generalization ability.
@@
We first collect a $\sim$1,000,000 sentence subsample of the Finnish mC4 corpus (Xue et al., 2021). We then extract unique words from the text and morphologically segment them using omorfi (Pirinen, 2015) and UralicNLP (Hämäläinen, 2019). After excluding words that analyzers did not cover, we manually annotate the segmentations to identify prefixes, lemmas, and affixes among the segments. We then perform stratified sampling based on the number of affixes to ensure an even range of morphological complexity in our data set. Finally, we extract sentences corresponding to each analyzed word from mC4 and validate whether they make sense. In a significant portion of cases, we notice that the raw sentences are noisy; in these cases, we opt to generate synthetic sentences using ChatGPT, which we (authors) then manually validate to be grammatical. Final data statistics and examples can be found in Appendix Tables 6 and 7.
@@ 4 Experiments
We treat the productivity task as an open-ended task in which the model is asked to derive a word from the given root and affixes and the systematicity task as a binary classification task in which the model is asked to determine whether the given derivation is grammatically correct. For the systematicity task, we generate negative examples by producing all the combinations of morphemes attached to the same root and choosing the top four compositions (two for morpheme lengths of 1 and 2) [^4] For 1-morpheme words, we manually annotate a negative morpheme to generate one negative option. that are closest to the original valid combination measured by the Levenshtein distance. We do this to ensure our incorrect combinations are challenging enough for the model as they will be deceptively close to a plausible derivation. We also experiment with other negative example selection strategies such as random selection and a heuristic selection based on the linguistic characteristics of the given language. We describe these settings in more detail and compare the results in Section 5.5. Finally, we (authors) manually verify all the generated negative examples and fix the label of false negatives.
@@
We evaluate several state-of-the-art multilingual instruction-finetuned LLMs, namely, two open-weights models, Aya-23 Aryabumi et al. (2024) and Qwen-2.5 Team (2024), and two closed-source models, Gemini-1.5 Gemini (2024) and GPT-4 OpenAI (2024). We evaluate all models on all languages except for Aya-23 which officially supports Turkish, but not Finnish. [^5] We also experimented with recent LLMs that are instruction-finetuned specifically on Finnish such as Poro-34B Luukkonen et al. (2024) and Ahma model series that are Llama Touvron et al. (2023) models fine-tuned on Finnish, however, we omitted them from our analysis as they failed to follow our task prompts in both English and Finnish templates.. We also report the performance of a random baseline that generates a derivation with a random combination of given morphemes (productivity task) and randomly decides whether the derivation is grammatically correct, and a majority baseline, which selects the most frequent label (in our case "No") for the systematicity task (not applicable for the productivity task). All models are evaluated using few-shot (1, 3, and 5) in-context learning and greedy decoding since our tasks are deterministic by nature [^6] We also experiment with other decoding strategies, however, find no significant difference in performance. Results for different decoding strategies can be found in Appendix A.5. Unless otherwise specified, prompt instructions are in English, and number of shots is set to $5$ for reported results [^7] We also experiment with paraphrased version of our prompt instructions, but find no significant difference in performance. Results for paraphrased prompt instruction can be found in Appendix A.6. Further details on model evaluation can be found in Appendix C.
@@
For the productivity task, we use Exact Match accuracy against the correct derivations. For the systematicity task, we report an average of Macro-F1 scores for each sample and a Coherence score that measures whether the model correctly and consistently identifies the validity (or invalidity) of all derivations for a given set of morphemes. Hence, coherence is defined as a binary score where the model gets a score of $1$ for a given sample if and only if it correctly guesses the validity of all derivations pertinent to that sample, otherwise 00. We employ this stringent metric to test the robustness of model performance similar to Storks and Chai (2021).
@@ 5.1 Effect of Morphological Complexity
Recent works have shown that morphological complexity plays a crucial role in the morphological generalization abilities of LLMs Anh et al. (2024); Czarnowska et al. (2019); Cotterell et al. (2018b). Morphological complexity is typically categorized into integrative (I-complexity) which refers to the predictability of inflected form and enumerative (E-complexity) complexity which refers to the number of cases and inflectional paradigms in language grammar Ackerman and Malouf (2013). While both languages we study are morphologically complex, our test suites include inflectional and derivational forms of varying length in the number of morphemes (1-7 in Turkish and 1-6 in Finnish). This allows us to study the effect of within-language E-complexity on the performance of our models. Figure 4 summarizes the GPT-4 performance for both tasks stratified by the number of bound morphemes on the Turkish data. On the productivity task, we observe a sharp downward trend (plummeting to nearly zero) in performance as the number of morphemes increases for both ID and OOD test suites with a relatively constant gap between ID and OOD performance while humans exhibit no such dependence on complexity (Appendix Tables 16, 25). This shows that humans learn their native language robustly and can easily produce and identify long novel words while models are quite sensitive to the morphological (E-) complexity.
@@ 5.3 Effect of Tokenization
Past work has shown that suboptimal tokenizers, especially byte-pair encoding Sennrich et al. (2016) used in GPT-4 have generally a negative effect on the morphological abilities of language models Meyer and Buys (2023); Bostrom and Durrett (2020); Hofmann et al. (2021). Whether the low performance of the model on the productivity task can be attributed to the suboptimal nature of the tokenization is of interest in particular because our tasks rely on the morphologically segmented morphemes while the model utilizes byte-level tokens that are mostly English. To measure the effect of the tokenization, we ran a version of the productivity task where the morphemes provided to the model are obtained by segmenting the final derivation based on the model’s own tokenizer instead of the morphologically-aligned units. Figure 6 compares the performance of the tokenizer-aligned morphemes with the morphologically-aligned morphemes on the ID test set. [^9] Since we use the word root as a definition for the nonce root and the tokenizer tends to break the words into meaningless chunks, we skip this experiment on the OOD test set. We see that the performance in both cases is very similar to each other which points to a possibility that tokenization may not be the underlying issue behind the low performance. This finding is also consistent with some past work on exploring morphological capabilities of ChatGPT Weissweiler et al. (2023) [^10] We note that we perform this analysis only with subword-level tokenizers, but not character-level tokenizers for two reasons: 1) To the best of our knowledge, at the time of writing this paper, there were no instruction-tuned multilingual language models for Turkish and Finnish that uses character-level tokenizers; 2) Past work has shown that character-level tokenizers do not offer any significant advantages over subword-level tokenizers in morphological generalization Libovický et al. (2021); Toraman et al. (2022).
@@ 6 Conclusion
In this paper, we proposed a novel experimental paradigm to test morphological generalization abilities of large language models through compositionality. Our tasks target measuring morphological productivity and systematicity in a given language. We applied these tasks on the morphologically complex languages of Turkish and Finnish and evaluated morphological compositional generalization abilities of several state-of-the-art large language models. Our experimental results and analysis reveal a significant gap in the performance of LLMs compared to humans with respect to generalization in morphology of agglutinative languages.
@@
While our novel tasks are language, dataset, and model-independent, our study only focused on two agglutinative languages and a few large language models. Therefore, the applicability of our findings in other languages and models should be further studied. We also mainly focused on the grammatical validity of the words, whereas it would be equally interesting to study the capacity of LLMs to produce and understand novel semantically and pragmatically valid derivations. While we have also optimized our prompts to be as simple and maximally instructive and tested in multiple languages and in chain-of-thought setting, whether a different set of prompts would produce the same results is not clear. Finally, we mainly evaluate models using greedy decoding due to the deterministic nature of our tasks and additionally only experiment with temperature and top-p sampling, however, the effect of different decoding strategies needs to be explored.